# Toward Scalable and Structured Global Station Weather Forecasting

### v1

@@ Abstract
Global Station Weather Forecasting (GSWF) is a key meteorological research area, critical to energy, aviation, and agriculture. Existing time series forecasting methods often ignore or unidirectionally model spatial correlation when conducting large-scale global station forecasting. This contradicts the intrinsic nature underlying observations of the global weather system, limiting forecast performance. To address this, we propose a novel Spatial Structured Attention Block in this paper. It partitions the spatial graph into a set of subgraphs and instantiates Intra-subgraph Attention to learn local spatial correlation within each subgraph, and aggregates nodes into subgraph representations for message passing among the subgraphs via Inter-subgraph Attention—considering both spatial proximity and global correlation. Building on this block, we develop a multiscale spatiotemporal forecasting model by progressively expanding subgraph scales. The resulting model is both scalable and able to produce structured spatial correlation, and meanwhile, it is easy to implement. The experimental results show that it can achieve performance improvements up to 16.8% over time series forecasting baselines at low running costs.
@@ 1 Introduction
To address the two limitations, we propose a novel Spatial Structured Attention Block that respects the first law of geography while also permitting long-distance message passing in this paper. The high-level idea is to partition the spatial graph into a set of subgraphs and instantiate self-attention to learn local spatial correlation within each subgraph (Intra-subgraph Attention). To capture the global spatial correlation, we further aggregate the nodes to produce subgraph representations and exchange information among the subgraphs via self-attention again (Inter-subgraph Attention). Since the entire block is differentiable, the message passed from distant nodes can be backpropagated through Inter-subgraph Attention to their correlated nodes in a parsimonious manner. To further enable the perception of spatial structure, we encode the shortest path distance between any two nodes as a bias term in the spatial attention mechanism. Moreover, we stack the proposed Spatial Structured Attention blocks with residual and gradually increase the subgraph scales to develop our eventual GSWF model. Such a design brings the following two appealing features: 1) our proposed method adopts the spatial graph to facilitate the perception and exchange of local spatial information, and thus it can yield sparse structure and reduce both the computational and memory burdens; 2) it permits message passing between distant node pairs in a parsimonious way, and thus it is capable of capturing global spatial correlation without incurring extra noise. To summarize:
@@
  • Building on the proposed block, we develop a multiscale GSWF model by gradually increasing the subgraph scales. The resulting model is scalable and can produce structured spatial correlation.
@@ 4.1 Spatial Information Preprocess
$\mathrm{SH}(\bm{\lambda,\phi})=\oplus_{n=0}^{N}\|_{l=0}^{\infty}\|_{m=-l}^{l}w^{m}_{l}{Y^{m}_{l}}(\bm{\lambda_{n},\phi_{n}})$ (2)
@@ 4.2 Spatial Structured Attention Block
$\bm{\alpha}_{p}=\mathrm{softmax}\left(\frac{\mathbf{Q}_{p}\mathbf{K}_{p}^{\top}}{\sqrt{d\,}}+\mathbf{SA^{bias}_{p}}\right)$ (11)
$\mathbf{SA^{bias}_{p}}=\sigma\left(\mathrm{SPD}\left(\mathbf{A}_{p}\right)\right)$ (12)
where $\mathbf{SA^{bias}_{p}}\in\mathbb{R}^{M\times M}$ is the intra-subgraph attention bias matrix in the $p$-th subgraph and $\sigma$ is an element-wise learnable scalar shared across all blocks. We can modify Eq. 7 in the same manner.
@@ 4.3 Multiscale Spatial Structured Architecture
Complexity Analysis. Our method achieves significant computational efficiency improvements by focusing on Intra-subgraph and Inter-subgraph Attention. Intra-subgraph Attention (for subgraphs with $\frac{N}{P}$ nodes) has complexity $\mathcal{O}(P(\frac{N}{P})^{2}D)=\mathcal{O}(\frac{N^{2}}{P}D)$, and Inter-subgraph Attention (for cross-subgraph correlation) has $\mathcal{O}(P^{2}D)$, leading to an overall complexity of $\mathcal{O}(\frac{N^{2}}{P}D+P^{2}D)$. Minimizing this (via derivative calculation) gives $\mathcal{O}(2N^{4/3}D)$ when $P=N^{2/3}$, with better scalability than quadratic-complexity methods. Our model’s efficiency can be further enhanced by linear attention; though Corrformer (specifically designed for GSWF tasks) has a better theoretical complexity of $\mathcal{O}(NT\log TD)$, it performs poorly in practice. This is due to its temporal alignment/rearrangement operations (for inferred node order), which disrupt data layout, increase memory latency (hindering hardware parallelism), and raise memory occupancy (via intermediate data storage). Detailed efficiency analysis is provided in Section 5.3.
@@ 6 Conclusion and Limitation
In this paper, we propose a novel Spatial Structured Attention Block that not only perceives spatial structure but also considers both spatial proximity and global correlation. Specifically, we partition the spatial graph into a set of subgraphs and utilize the Intra-subgraph and Inter-subgraph Attention to learn local and global spatial correlation. Building on the proposed block, we develop a multiscale model by progressively increasing the subgraph scales. The resulting model is scalable and can produce structured spatial correlation. Performance comparison and efficiency analysis validate the superiority of our method in medium and large-scale global station weather forecasting scenarios. Ablation studies confirm the effectiveness of the model designs, and we further explore the hyperparameters in model construction.

### v2

# S2Transformer: Scalable Structured Transformers for Global Station Weather Forecasting

@@ Abstract
Global Station Weather Forecasting (GSWF) is a key meteorological research area, critical to energy, aviation, and agriculture. Existing time series forecasting methods often ignore or unidirectionally model spatial correlation when conducting large-scale global station forecasting. This contradicts the intrinsic nature underlying observations of the global weather system, limiting forecast performance. To address this, we propose a novel Spatial Structured Attention Block in this paper. It partitions the spatial graph into a set of subgraphs and instantiates Intra-subgraph Attention to learn local spatial correlation within each subgraph, and aggregates nodes into subgraph representations for message passing among the subgraphs via Inter-subgraph Attention—considering both spatial proximity and global correlation. Building on this block, we develop a multiscale spatiotemporal forecasting model S2Transformer by progressively expanding subgraph scales. The resulting model is both scalable and able to produce structured spatial correlation, and meanwhile, it is easy to implement. The experimental results show that it can achieve performance improvements up to 16.8% over time series forecasting baselines at low running costs.
@@ 1 Introduction
To address the two limitations, we propose a novel Spatial Structured Attention Block that respects the first law of geography while also permitting long-distance message passing in this paper. The high-level idea is to partition the spatial graph into a set of subgraphs and instantiate self-attention to learn local spatial correlation within each subgraph (Intra-subgraph Attention). To capture the global spatial correlation, we further aggregate the nodes to produce subgraph representations and exchange information among the subgraphs via self-attention again (Inter-subgraph Attention). Since the entire block is differentiable, the message passed from distant nodes can be backpropagated through Inter-subgraph Attention to their correlated nodes in a parsimonious manner. To further enable the perception of spatial structure, we encode the shortest path distance between any two nodes as a bias term in the spatial attention mechanism. Moreover, we stack the proposed Spatial Structured Attention blocks with residual and gradually increase the subgraph scales to develop our eventual GSWF model S2Transformer. Such a design brings the following two appealing features: 1) our proposed method adopts the spatial graph to facilitate the perception and exchange of local spatial information, and thus it can yield sparse structure and reduce both the computational and memory burdens; 2) it permits message passing between distant node pairs in a parsimonious way, and thus it is capable of capturing global spatial correlation without incurring extra noise. To summarize:
@@
  • Building on the proposed block, we develop a multiscale GSWF model S2Transformer by gradually increasing the subgraph scales. The resulting model is scalable and can produce structured spatial correlation.
@@ 4.1 Spatial Information Preprocess
$\mathrm{SH}(\bm{\lambda,\phi})=\oplus_{n=0}^{N}\|_{l=0}^{\infty}\|_{m=-l}^{l}w^{m}_{l}{Y^{m}_{l}}({\lambda_{n},\phi_{n}})$ (2)
@@ 4.2 Spatial Structured Attention Block
$\bm{\alpha}_{p}=\mathrm{softmax}\left(\frac{\mathbf{Q}_{p}\mathbf{K}_{p}^{\top}}{\sqrt{d\,}}+\mathbf{SA}^{\mathrm{bias}}_{p}\right)$ (11)
$\mathbf{SA}^{\mathrm{bias}}_{p}=\sigma\left(\mathrm{SPD}\left(\mathbf{A}_{p}\right)\right)$ (12)
where $\mathbf{SA}^{\mathrm{bias}}_{p}\in\mathbb{R}^{M\times M}$ is the intra-subgraph attention bias matrix in the $p$-th subgraph and $\sigma$ is an element-wise learnable scalar shared across all blocks. We can modify Eq. 7 in the same manner.
@@ 4.3 Multiscale Spatial Structured Architecture
[…]
@@ 6 Conclusion and Limitation
In this paper, we propose a novel Spatial Structured Attention Block that not only perceives spatial structure but also considers both spatial proximity and global correlation. Specifically, we partition the spatial graph into a set of subgraphs and utilize the Intra-subgraph and Inter-subgraph Attention to learn local and global spatial correlation. Building on the proposed block, we develop a multiscale model S2Transformer by progressively increasing the subgraph scales. The resulting model is scalable and can produce structured spatial correlation. Performance comparison and efficiency analysis validate the superiority of our method in medium and large-scale global station weather forecasting scenarios. Ablation studies confirm the effectiveness of the model designs, and we further explore the hyperparameters in model construction.